# Kang M, Chen D, Pu Z, Gao J, Yu W. Spatio-Temporal Graphical Counterfactuals: An Overview

### v2

@@ Abstract
Counterfactual thinking is a critical yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve their performances for new scenarios. Many research works, including Potential Outcome Model and Structural Causal Model, have been proposed to realize it. However, their modelings, theoretical foundations and application approaches are usually different. Moreover, there is a lack of graphical approach to infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between multiple units. Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal framework to infer the spatio-temporal counterfactuals.
@@ 1 Introduction
How to enable an intelligent machine to think and answer a spatio-temporal counterfactual question? E.g., sometimes we want to know, if a different investment strategy had been implemented, would we obtain higher returns? Or, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? Or, would someone still purchase the product even if they had never been shown the advertisement? These questions generally have the spatio-temporal counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”. Thus, according to the pattern, firstly, time is necessary to consider, because the questions are queried now, but the counterfactual actions are supposed to be done in the past. Secondly, counterfactual outcomes are unobservable. The outcomes of some actions are observed if the actions have been done to the real-world system, thus these outcomes are factual. But the counterfactual outcomes are not observed, and they are imaged through supposing different actions had been done. Thirdly, the real-world system can be observed as multivariate time series, and there are widespread multivariate interactions within the time series [1, 2].
@@
To enable an intelligent machine to imitate counterfactual thinking like human, thus, Potential Outcome Model (POM) [3] and Structural Causal Model (SCM) [4, 5, 6, 7] are proposed as two frameworks with different foundations. POM uses stable unit treatment value assumption (SUTVA), consistency assumption, positivity assumption and ignorability assumption as foundations [8, 9, 10, 11]. If these assumptions are satisfied, one can use matching or imputation approaches to infer counterfactuals without confounders. But in contrast, SCM builds a causal ladder, i.e., abduction, action and prediction, to infer counterfactuals through graphical language d-separation [4, 5, 7]. With the graphical language d-separation, then causal Markov assumption, faithfulness assumption and causal sufficiency assumption are proposed to guarantee the correctness of counterfactual inference on SCM [12, 13, 14, 15, 16].
@@
However, the different foundations between POM and SCM lead to differences in counterfactual inference. Such differences include: (i) POM assumes the positivity, but SCM does not. That means the counterfactuals in POM are factual actually and can be found in observations certainly. But in contrast, the counterfactuals in SCM can be out of observations and unfalsifiable, that is, one cannot validate the correctness of the inferred counterfactuals. (ii) POM assumes the consistency, but it is rejected for SCM [7]. Actually, SCM also needs the consistency assumption implicitly to guarantee that the counterfactual outcome is one of the unobserved potential outcomes. (iii) SCM uses a type of graphical language to deconfound with certain rules, but POM mainly relies on human-guided deconfounding, that makes it difficult to satisfy the ignorability. But in contrast, SCM can deconfound through back-door criterion [17, 18, 4, 5] based on the casual sufficiency assumption, and it is equivalent to the ignorability actually. (iv) It is difficult for POM to analyze indirect causation (mediation), thus POM typically relies on Baron-Kenny model [19, 20] for assistance. By contrast, it is easier for SCM by using front-door criterion [17, 18, 4, 5], that decomposes a causal pathway into two parts, i.e., from source variable to mediator and from mediator to target variable, and then conduct back-door adjustment twice if no other mediator.
@@
Moreover, whether it is POM or SCM currently, they are both unable to model the spatio-temporal interactions between system units. E.g., the SUTVA requires no interaction between system units in POM. Moreover, SCM based on the traditional Bayesian network has no modeling for multivariate time series, thus the nonstationarity in time series is often neglected. Thus, this motivates us to find a common point to unify the two frameworks, and strengthen the inference for the spatio-temporal counterfactuals. Fortunately, our recent work [21, 22] has proposed a Spatio-Temporal Bayesian Networks (STBNs) model based on temporal assumption, that assumes the cause precedes or parallels the effect in time. And it is found that, the famous complex networks [23, 24, 25], a type of network dynamics to model spatio-temporal interactions, are a special case of STBNs.
@@
Thus, we further conduct this work to build a spatio-temporal graphical counterfactual framework based on STBNs. The framework requires two components: (i) an STBNs used as the foundation of SCM, (ii) a Forward Counterfactual Inference Algorithm designed on the decomposability of the front-door criterion. Then we discuss how to get the two points, following the storyline as shown in Fig. 1. Thus, the main contributions of this work are as follows:
@@
    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section 2 to Section 4.
@@
    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss Spatio-Temporal Bayesian Networks, its nonstationarity, and the relationship with complex networks. This part is organized in Section 5.
@@ 2 Potential Outcome Model
The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in Table 1. Here, $Y$ is the potential outcomes of the target variable. $T$ is treatment variable. $T=1$ represents being treated and $T=0$ represents being controlled, with a group of randomized control trials on $N$ units. $X$ represents covariates that have relationship with $Y$ and $T$, and are controlled for randomized trials. POM models the general process from causal estimands ($T$, $Y$ and $X$) to “Science” [26]. In this case, for different treatments $T$, only one potential outcome, $Y(1)$ or $Y(0)$, can be observed, and another one is unobserved, namely the counterfactual.
@@ 2.1 Inferring Counterfactuals through POM
To infer the counterfactuals through POM, one common approach is matching, e.g., exact matching. For a unit $i$, one can search the samples with exactly matching covariates $X$ but treatment $T$ is different, and then estimate the counterfactual outcome $Y_{i}(1-T_{i})$. But it is difficult to conduct if the dimension of $X$ is high, that requires large number of samples to support the matching. Thus, approximate matching is commonly better, e.g., caliper matching [27, 28, 29, 30] and propensity score matching [11, 31, 32, 33]. They approximately search a group of similar samples for the target unit with one or multiple measurements, and finally calculate the counterfactuals.
Another common approach is data-driven imputation, that views the counterfactuals as missing values, and interpolates them by fitting the observational data. To realize it, linear or nonlinear regressions [34, 35, 36, 37] are commonly used to interpolate the missing values from the trained regressive models on the observational data. Different from this, tensor decomposition approaches [38, 39, 40] recover the missing values by decomposing and reconstructing the sparse data tensor. Moreover, deep generative models [41, 42, 43, 44, 45, 46, 47] are also used to model the uncertainty in data, and generate the missing values by randomly sampling. Note that, the generated missing values are not unique, but a group of values conforming to a random distribution. This is different compared to the regressive approaches and tensor decomposition approaches.
@@ 2.2 Foundational Assumptions of POM
Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is SUTVA, which is defined as {assumption}[SUTVA [10]] In Table 1,
@@
    there is no hidden versions of treatments, that is, no matter how unit $i$ received treatment $T=1$, the outcome that would be observed would be $Y_{i}(1)$, and similarly for treatment $T=0$.
SUTVA guarantees the identical independence between the experimental units, and for unit $i$, the outcome $Y_{i}$ is only up to its treatment $T_{i}$, not the others’. However, this assumption is not always satisfied, e.g., in the case of social network, in which there are frequent interactions between units with varying time.
@@
where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in Table 1. This is the same to the case of conditional expectation with $X$. Thus, the consistency assumption guarantees the missing counterfactual values in Section 2.1 can be interpolated by the other observational values actually.
The third assumption is positivity, defined as {assumption}[Positivity [11]] In Table 1, $0<P(T=1|X=x)<1$. This means for $X=x$, there are always some treatments randomly assigned with $T=0$ and $T=1$. Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same to the counterfactuals for $T=0$.
The forth assumption is ignorability, defined as {assumption}[Ignorability [8, 9]] In Table 1, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit. Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere the observations of $Y(1),Y(0)$. Thus, the ignorability assumption is also called as unconfoundedness. This assumption are usually made to avoid the preference on experimental manipulation, that would guarantee the randomization in the experiments [48].
@@ 3 Structural Causal Model
Different from POM, SCM provides a type of graphical language to discuss causality intuitively, which avoids complex human-guided technical operations. An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$. Note that, endogenous variables are observable from data, but exogenous variables are not. Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ are also set to describe the functional relationship between these variables.
@@
Note that, in an SCM, the functional relationship must conform to a directed-acyclic graphical constraint. The graph is early originated from the concept of Bayesian network [12, 13, 14, 15, 16]. And latter, it is also called as causal graphical model to highlight the causality [49, 17, 18, 4, 5]. And recently, it is also called as causal network in the field of causal discovery [50, 51, 52, 53, 54]. But actually, they are the same in the framework of SCM, and they are all directed acyclic graph (DAG) that represents Markovian knowledge. Thus, to avoid misunderstanding, we here use the name of Bayesian network uniformly, due to the fact that causal graphical model and causal network are not always a Bayesian network in Fig. 2. For example, Direct Cyclic Graph [55, 56, 57, 58], Markov network [59, 60], and Full Time Graph [6], they are also causal graph or causal network.
Figure 2: An example of SCM [5].
@@ 3.1 Pearl’s Causal Ladder [4, 5, 7]
Figure 3: A diagram of Pearl’s causal ladder [5].
As shown in Fig. 3, we query “observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”. Obviously, this is a counterfactual question. Thus, follow the Pearl’s causal ladder, as shown in Fig. 3, we first infer the values of exogenous variables $\mathbf{U}=\{U_{1},U_{2},U_{3}\}$ from the observational data $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, as follows:
@@
The intervention, also named as $do$-operator [4], can be generalized into the form of probability, if modularity assumption is introduced as follow: {assumption}[Modularity [4]] If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
@@
    if $X\notin\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X|\mathbf{Pa}(X))$ remains unchanged. Here, $\mathbf{Pa}(X)$ is the predecessors of $X$ in Bayesian network, that is also called as causal parents.
@@
This means, if a group of interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$, the values of them would be fixed. The connections between them and their causal parents would be broken, and their causal parents would not affect them anymore in the modified SCM, just like in Fig. 3. Moreover, this also means that the probability distributions of the other variables that are not be intervened would not change.
@@ 3.2 Foundational Assumptions of Bayesian Network
As presented above, to clearly define the causality, every SCM is associate to a Bayesian network with directed-acyclic graphical constraint. Thus, the causality in Bayesian network determines the counterfactuals foundationally. Thus, we discuss the foundational assumptions of Bayesian network in the following contents.
@@
for all disjoint subsets $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$. It is intuitive to understand the two assumptions, that is, the Bayesian network $\mathcal{G}$ has one-to-one correspondence to the independence of probability distribution $P$ in observational data. Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follow:
@@
Moreover, a Bayesian network is also assumed to be causally sufficient, that is {assumption}[Causal Sufficiency [12, 13, 14, 15, 16]] Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$. This means, if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network. Actually, the ignorability assumption in POM (see Assumption 2.2) also implies the causal sufficiency, and this assumption is weaker than ignorability. But note that, the causal sufficiency is hard to satisfy practically, because many real-world systems (e.g., economic system and climate system) are complex, thus, we could not collect all the information to describe the systems usually. Thus, some research works focus on this issue. For example, FCI (or Fast Causal Inference) algorithm and its variants [61, 62, 63, 64, 65, 66] are proposed to discover causality in the presence of hidden variables. And some other related works [67, 68, 69] investigate the causal discovery with soft intervention, that intervenes the SCM but does not change the network structure. There are also some related works [70, 71, 72, 73] proposed approaches to search the optimal and efficient adjustment set, that all its variables are observable, minimal and valid, and the removal of any of its variables would destroys the validity.

### v3

# Kang M Y, Chen D X, Pu Z Y, et al. Spatio-Temporal Graphical Counterfactuals: An Overview

@@ Abstract
Counterfactual thinking is a crucial yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve performance for new scenarios. Many research works, including the Potential Outcome Model (POM) and the Structural Causal Model (SCM), have been proposed to address this. However, their modeling, theoretical foundations, and application approaches often differ. Moreover, there is a lack of graphical approaches for inferring spatio-temporal counterfactuals, that account for spatial and temporal interactions among multiple units. Thus, in this work, we aim to present a survey that compares and discusses different counterfactual models, theories and approaches. Additionally, we propose a unified graphical causal framework to infer spatio-temporal counterfactuals.
@@ 1 Introduction
How can we enable an intelligent machine to think about and answer a spatio-temporal counterfactual question? For example, sometimes we want to know: if a different investment strategy had been implemented, would we have obtained higher returns? Or, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? Or would someone still purchase a product even if they had never seen the advertisement? These questions generally follow a spatio-temporal counterfactual pattern: “Given the observed $\dots$, if $\dots$ had been done, how would $\dots$ have evolved?”. Thus, according to the pattern, time must be considered first, because the questions are posed now, but the counterfactual actions are supposed to be performed in the past. Second, counterfactual outcomes are unobservable. The outcomes of some actions are observed when they are performed on a real-world system, thus, these outcomes are factual. But the counterfactual outcomes are not observed, and they are imagined by supposing that different actions had been taken. Third, the real-world system can be observed as a multivariate time series, and there are widespread multivariate interactions within it [1, 2].
@@
To enable intelligent machines to mimic counterfactual thinking, the Potential Outcome Model (POM) [3] and Structural Causal Model (SCM) [4, 5, 6, 7] have been proposed as two frameworks with different foundations. POM uses stable unit treatment value assumption (SUTVA), consistency assumption, positivity assumption and ignorability assumption as foundations [8, 9, 10, 11]. If these assumptions are hold, one can use matching or imputation approaches to infer counterfactuals without confounders. In contrast, SCM builds a causal ladder, i.e., abduction, action and prediction, to infer counterfactuals through graphical language d-separation [4, 5, 7]. With the graphical language d-separation, the causal Markov assumption, faithfulness assumption and causal sufficiency assumption are proposed to guarantee the correctness of counterfactual inference on SCM [12, 13, 14, 15, 16].
@@
However, the distinct foundations between POM and SCM lead to differences in counterfactual inference. These differences include: (i) POM assumes positivity, whereas SCM does not. Thus, the counterfactuals in POM are factual and can be found in observations certainly. In contrast, counterfactuals in SCM can be out of observations and are therefore unfalsifiable, that is, their correctness cannot be empirically validated. (ii) POM explicitly assumes consistency, but this assumption is rejected for SCM [7]. Actually, SCM also implicitly relies on the consistency assumption to ensure that the counterfactual outcome corresponds to one of the possibly unobserved potential outcomes. (iii) SCM uses a type of graphical language to deconfound with certain rules, but POM mainly relies on human-guided deconfounding, which makes it difficult to satisfy the ignorability. In contrast, SCM can deconfound via the back-door criterion [17, 18, 4, 5] under the casual sufficiency assumption, and it is equivalent to ignorability. (iv) POM faces limitations to analyze indirect causation (mediation), thus POM typically relies on the Baron-Kenny model [19, 20] for assistance. By contrast, SCM facilitates this analysis through the front-door criterion [17, 18, 4, 5], which decomposes a causal pathway into two segments: from the source variable to mediator, and from the mediator to target variable. Back-door adjustment is then applied twice, assuming no additional mediators are presented.
@@
Moreover, both POM and SCM currently lack the ability to model the spatio-temporal interactions among system units. For example, the SUTVA requires no interaction between system units in POM. Similarly, SCM, when based on the traditional Bayesian network, does not handle multivariate time series and often overlooks nonstationarity. This limitation motivates us to identify a common point to unify the two frameworks and strengthen inference for spatio-temporal counterfactuals. Our recent work [21, 22] has proposed a spatio-temporal Bayesian networks (STBNs) model based on a temporal assumption that the cause precedes or parallels the effect in time. It is found that the famous complex networks [23, 24, 25], a type of network dynamics used to model spatio-temporal interactions, are also a special case of STBNs.
@@
Thus, we further conduct this work to build a spatio-temporal graphical counterfactual framework based on STBNs. This framework requires two components: (i) STBNs used as the foundation of SCM, (ii) a Forward Counterfactual Inference Algorithm designed on the decomposability of the front-door criterion. We then discuss how to get the two points, following the storyline as shown in Fig. 1. Thus, the main contributions of this work are as follows:
@@
    A survey is conducted to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section 2 to Section 4.
@@
    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss the spatio-temporal Bayesian networks, their nonstationarity, and the relationship with complex networks. This part is organized in Section 5.
@@ 2 Potential Outcome Model
The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in Table 1. Let $Y$ be the potential outcomes of the target variable, and $T$ is the treatment variable. $T=1$ represents assignment to the treatment group, and $T=0$ represents assignment to the control group in a randomized controlled trial involving $N$ units. $X$ represents covariates that have relationship with $Y$ and $T$, and are controlled for randomized trials. POM models the general process from causal estimands ($T$, $Y$ and $X$) to “Science” [26]. In this case, for different treatments $T$, only one potential outcome, $Y(1)$ or $Y(0)$, can be observed, and another one is unobserved, namely the counterfactual.
@@ 2.1 Inferring Counterfactuals through POM
To infer the counterfactuals using POM, one common approach is exact sample matching. For a unit $i$, one can search the samples with exactly matching covariates $X$ but treatment $T$ is different, and then estimate the counterfactual outcome $Y_{i}(1-T_{i})$. But it is difficult to conduct when the dimension of $X$ is high, as it requires a large number of samples to support the matching. Thus, approximate matching is commonly better, e.g., caliper matching [27, 28, 29, 30] and propensity score matching [11, 31, 32, 33]. They search for a group of similar samples to the target unit with one or more measurements, then calculate counterfactuals.
Another common approach is data-driven imputation, which views counterfactuals as missing values, and interpolates them by fitting the observational data. Linear or nonlinear regressions [34, 35, 36, 37] are typically used to interpolate missing values based on patterns learned from the observational data. In contrast, tensor decomposition approaches [38, 39, 40] recover the missing values by decomposing and reconstructing a sparse data tensor. Deep generative models [41, 42, 43, 44, 45, 46, 47] have also been used to capture data uncertainty and generate missing values by randomly sampling. Note that, the imputed missing values produced by generative models are not unique. Instead, they constitute a distribution of plausible values, which distinguishes them from regression-based and tensor decomposition approaches.
@@ 2.2 Foundational Assumptions of POM
Before using the above approaches within POM framework, four assumptions must be accepted. The first is SUTVA, which is defined as {assumption}[SUTVA [10]] In Table 1,
@@
    There are no hidden versions of treatments, that is, no matter how unit $i$ received treatment $T=1$, the outcome that would be observed would be $Y_{i}(1)$, and similarly for treatment $T=0$.
SUTVA guarantees the identical independence between the experimental units, and for unit $i$, the outcome $Y_{i}$ is only up to its treatment $T_{i}$, not the others’. However, this assumption is not always satisfied, e.g., in social networks, where interactions occur at varying times.
@@
where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in Table 1. This is the same as the case of taking a conditional expectation with respect to $X$. Thus, the consistency assumption guarantees that the missing counterfactual values in Section 2.1 can be interpolated by the other observational values.
The third assumption is positivity, defined as {assumption}[Positivity [11]] In Table 1, $0<P(T=1|X=x)<1$. This means that, for $X=x$, some treatments are always assigned to $T=0$ or $T=1$ at random. Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same as the counterfactuals for $T=0$.
The fourth assumption is ignorability, defined as {assumption}[Ignorability [8, 9]] In Table 1, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit. Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere with the observations of $Y(1),Y(0)$. Thus, the ignorability assumption is also called unconfoundedness assumption. This assumption is usually made to avoid the preference for experimental manipulation, which would guarantee randomization in the experiments [48].
@@ 3 Structural Causal Model
In contrast to POM, SCM provides a graphical language for intuitively discussing causality, thereby avoiding complex human-guided technical operations. An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$. Note that, endogenous variables are observable from data, but exogenous variables are not. Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ can also be used to describe the functional relationship between these variables.
Figure 2: Example of SCM [5].
Figure 3: Diagram of Pearl’s causal ladder [5].
@@
In an SCM, the functional relationship must conform to the directed acyclic graphical constraint. The graph is originally based on the concept of Bayesian network [12, 13, 14, 15, 16]. And later, it is also called as causal graphical model to highlight its causal structure [49, 17, 18, 4, 5]. And recently, it is also called as causal network in the field of causal discovery [50, 51, 52, 53, 54]. But actually, they are the same in the framework of SCM, and they are all directed acyclic graph (DAG) that represents Markovian knowledge. Thus, to avoid misunderstanding, we here use the name of Bayesian network uniformly, due to the fact that causal graphical model and causal network are not always a Bayesian network in Fig. 3. For example, direct cyclic graphs [55, 56, 57, 58], Markov networks [59, 60], and full time graph [6], they are also causal graph or causal network.
@@ 3.1 Pearl’s Causal Ladder [4, 5, 7]
As shown in Fig. 3, consider the following query “observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”. Obviously, this is a counterfactual question. Thus, following Pearl’s causal ladder, as shown in Fig. 3, we first infer the values of exogenous variables $\mathbf{U}=\{U_{1},U_{2},U_{3}\}$ from the observational data $X_{1}=0.5,X_{2}=1$ and $X_{3}=1.5$, as follows:
@@
The intervention, also named as $do$-operator [4], can be generalized into the form of probability, if the modularity assumption is introduced as follows: {assumption}[Modularity [4]] If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
@@
    if $X\notin\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X|\mathbf{Pa}(X))$ remains unchanged. Here, $\mathbf{Pa}(X)$ denotes the set of predecessors of $X$ in a Bayesian network, also called the causal parents of $X$.
@@
This means that if a set of interventions is applied to $X_{j_{1}},\dots,X_{j_{p}}$, their values would be fixed. The connections between them and their causal parents would be broken, and their causal parents would not affect them anymore in the modified SCM, just like in Fig. 3. Moreover, this also means that the probability distributions of the other variables that are not intervened upon would not change.
@@ 3.2 Foundational Assumptions of Bayesian Network
As presented above, to clearly define causality, each SCM is associated with a Bayesian network satisfying the directed acyclic graphical constraint. Thus, the causality in the Bayesian network provides the foundation for counterfactuals. Thus, we discuss the foundational assumptions of Bayesian networks in the following sections.
@@
for all disjoint subsets $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$. It is intuitive to understand the two assumptions: the Bayesian network $\mathcal{G}$ corresponds one-to-one to the independence structure of the probability distribution $P$ in the observational data. Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follows:
@@
Moreover, a Bayesian network is also assumed to be causally sufficient, that is {assumption}[Causal Sufficiency [12, 13, 14, 15, 16]] Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$. This means that if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network. Actually, the ignorability assumption in POM (see Assumption 2.2) also implies the causal sufficiency, and this assumption is weaker than ignorability. But note that, causal sufficiency is hard to satisfy in practice, because many real-world systems (e.g., the economic and climate systems) are complex, thus, we usually cannot collect all the information needed to describe them. Thus, some research works focus on this issue. For example, the Fast Causal Inference algorithm and its variants [61, 62, 63, 64, 65, 66] are proposed to discover causality in the presence of hidden variables. And some other related works [67, 68, 69] investigate causal discovery with soft intervention, which intervenes in the SCM without altering the network structure. Related works [70, 71, 72, 73] have also proposed approaches for identifying optimal and efficient adjustment sets, in witch all variables are observable, minimal and valid, and or which removing any variables would invalidate the adjustment.